OpenStudy (jenniferjuice):

Choose the correct x-intercept and y-intercept of the equation 3x − 2y = 6. x-intercept: (0, −3), y-intercept: (0, 2) x-intercept: (−3, 0), y-intercept: (2, 0) x-intercept: (0, 2), y-intercept: (0, −3) x-intercept: (2, 0), y-intercept: (0, −3)

OpenStudy (anonymous):

for solving x-intercept put y=0 3x-2(0)=6 3x-0=6 3x=6 x=6/3 x=2 hence the point be (2,0) and for solving y-intercept put x=0 3(0)-2y=6 -2y=6 2y=-6 y=-6/2 y=-3 hence the point (0,-3) 4th option was correct